# AniSora: Exploring the Frontiers of Animation Video Generation in the Sora Era

### v3

@@ 5 Experiment
Table 1: Automated Performance Comparison of Different Methods. (Note that AniSora-K denotes the results with keyframe interpolation, and AniSora-I denotes the interpolated average results of AniSora)
Method Appearance Consistency
Motion Motion Aesthetic Imaging I2V I2V Overall Subject
Smoothness Score Quality Quality Subject Background Consistency Consistency
Opensora-Plan(V1.3) 99.13 76.45 53.21 65.11 93.53 94.71 21.67 88.86
Opensora(V1.2) 98.78 73.62 54.30 68.44 93.15 91.09 22.68 87.71
Vidu 97.71 77.51 53.68 69.23 92.25 93.06 20.87 88.27
Cogvideo(5B-V1) 97.67 71.47 54.87 68.16 90.68 91.79 21.87 90.29
MiniMax 99.20 66.53 54.66 71.67 95.95 95.42 21.82 93.62
AniSora 99.34 45.59 54.31 70.58 97.52 95.04 21.15 96.99
AniSora-K 99.12 59.49 53.76 68.68 95.13 93.36 21.13 94.61
AniSora-I 99.31 54.96 54.67 68.98 94.16 92.38 20.47 95.75
GT 98.72 56.05 52.70 70.50 96.02 95.03 21.29 94.37
@@ 5.1 Benchmark Evaluation
In this section, we give both objective and human evaluation results of our benchmark.
Automated Evaluation. To obtain the objective results, we choose several dimensions in VBench [11], e.g., motion smoothness, aesthetic quality, imaging quality, subject consistency, I2V subject consistency, I2V background consistency, and overall consistency. The former three metrics evaluate the visual quality, while the latter four reflect the degree of consistency. Especially, in VBench, overall consistency evaluates the text-video consistency, since they use ViCLIP [26] as the baseline model. In addition, we utilize a motion amplitude model, which is based on ActionCLIP [24] framework to evaluate the motion score of the generation clips. In detail, About $10$ million animation video clips and their corresponding motion captions are collected into $6$ degrees of movement amplitude (from stillness to significant motion) to finetune the action model. Finally, the motion score is obtained from the similarity score between the designed motion prompt and the participant video.
$S_{motion}=Cos(MCLIP(V),MCLIP(T_{m})),$ (2)
where $MCLIP$ denotes the finetuning action model. $V$ represents the generation video, and $T_{m}$ denotes the designed motion prompt.
$6$ recent I2V investigations are involved into our evaluation: Open-sora-V1.2 [32], Open-sora-plan-V1.3 [14], Cogvideox-5B-V1 [29], Vidu [23], Minimax [16] and AniSora(ours).
Tab. 1 gives the automated results from $8$ metrics. We observe that our method performs well on subject consistency and motion smoothness, and closely on other $5$ dimensions except motion score. These mainly because we conduct a thorough assessment of the balance between generation quality and motion magnitude, and find most generation clips with big motion results in distortion or unnatural segments. It is worth mentioning that the automated scores of AniSora are similar to those of GT, and the visual performance can refer to Fig. 2. Furthermore, the automated scores from VBench show the room for improvement across several dimensions, and we will provide our improved metrics soon.
Table 2: Human Evaluation Scores
Method Appearance Consistency
Visual Visual Visual Text-Video Image-Video Character
Smooth Motion Appeal Consistency Consistency Consistency
Opensora-Plan(V1.3) 38.1 38.92 47.88 55.82 43.52 34.72
Opensora(V1.2) 28.14 37.24 37.46 47.64 42.62 31.52
Vidu 58.78 47.9 65.48 60.8 56.5 54.26
CogVideoX(5B-V1) 46.64 49.3 56.06 68.82 56.36 48.88
MiniMax 65.98 57.08 71.56 80.38 67.88 65.82
AniSora(Ours) 71.68 51.06 74.36 71.56 78.38 75.14
Human Evaluation. To comprehensively evaluate our model, we introduce a brief and clear human blind testing for $6$ dimensions: visual smoothness, visual motion, visual appeal, text-video, image-video, and character consistency. Correspondingly, visual smoothness, text-video, and character consistency are similar to motion smoothness, overall, and subject consistency, respectively. Moreover, image-video consistency is equal to I2V subject and I2V background consistency, visual appeal is equal to aesthetic quality and imaging quality, and visual motion is the same as the motion score mentioned in the automated evaluation. In detail, each participant labeled $6$ dimensions (from $1$ to $5$, and $5$ is the best) without prior knowledge of the generation methods. Tab. 2 shows the human evaluation results in a percentage format. We observe that Anisora outperforms the other methods across most dimensions; however, there is still substantial room for improvement, particularly in text-video consistency. We conducted a statistical analysis to evaluate the consistency of scores given by 12 raters across various dimensions. The results indicate that the Pearson correlation coefficients for individual dimensions range from 0.5 to 0.6, with an overall correlation coefficient of 0.56. This suggests that even human evaluators exhibit significant subjectivity and randomness when assessing the quality of generated videos across different dimensions. These findings highlight the importance of establishing consistent and objective evaluation criteria for assessing video generation quality.
@@ 5.2 Spatiotemporal Mask module
Frame Interpolation. Tab. 1 presents the results of different interpolation settings on our benchmark dataset (AniSora-K and AniSora-I). Our evaluation process involved generating videos on our benchmark with various guidance conditions sampled at equal proportions, which can refer to Fig. 3. We then compute the average score of all samples, as well as specific statistical analysis for keyframe interpolation results. The performance indicates that single-frame guidance achieves competitive results whether the guiding frame is placed at the beginning, middle, or end of the frame sequence, which also consistently outperforms other methods. Adding more guiding frames further improves both character consistency and motion stability. We also observed from the motion score and smooth score that our baseline model achieves a balance between motion range and consistency, while keyframe guidance enables the model to produce animation videos with larger motion ranges and more realistic motion. More samples can be found in supplementary materials.
Motion Area Condition. The evaluation of motion area condition is constructed based on our benchmark dataset. For each initial frame, we performed saliency segmentation, followed by connected-component analysis to generate bounding boxes for each instance. Then we manually filtered the results to select high-quality motion area masks, resulting in 200 samples. Following the experiment settings in [7], we conducted the comparison of motion mask precision in Tab. 3. We also computed the score of AnimateAnything on our selected 200 samples. The lower score is primarily due to flickering and noise appearing outside the motion mask area. The results demonstrate the effectiveness of our spatial mask module in controlling movable regions. It is also noticeable that even without motion control, our generation model trained for animation video still shows a certain level of control. This may be due to the effective prompt-based guidance for the main subject, which aligns well with the defined motion mask. Fig. 5 also illustrates several motion mask guidance examples.
@@
5.3 Animation Video Training
Figure 6: Comparison of 2D and 3D animation examples. The badcases in 2D animation are mainly due to exaggerated deformations, diverse appearances, and motions that violate physical laws.
Table 4: Human Evaluation Results between 2D and 3D Generation Clips
Dims 2D 3D All
Visual Smooth 70.23 73.48 71.68
Visual Motion 51.14 50.97 51.06
Visual Appeal 74.05 74.74 74.36
Text-Video Consistency 70.23 73.21 71.56
Image-Video Consistency 77.59 79.37 78.38
Character Consistency 75.64 74.52 75.14
Figure 7: Comparison of results w/wo multi-task learning. The highlighted regions in red demonstrate significant improvements in stability and consistency after applying multi-task learning.
2D and 3D Animation. Analysis using QWEN2 [25] shows that 2D samples account for 85% of our dataset, yet 3D animation generation quality consistently surpasses that of 2D. Benchmark evaluations in Tab. 4 confirm 3D animations demonstrate superior visual appearance and consistency, a phenomenon unique to animation training. We attribute this gap to the pre-trained model’s exposure to real-world video data. Unlike 2D animations with diverse motion patterns, 3D animations rendered by physics-based engines like Unreal Engine follow consistent physical laws, enabling better knowledge transfer during SFT. Consequently, improving generalization on 2D animation data remains more challenging than on 3D or real-world data.
The Fig. 6 demonstrates some results of 2D and 3D animation generation. Artifacts are more prevalent in 2D generation results, such as exaggerated deformations, more diverse character appearances, and motions that break the physical rules. For instance, in the third row of 2D examples, tears appear to be floating in the air, making it more difficult for the model to capture the dynamic details accurately. In contrast, the motions rendered by the physics-based engines in 3D animations enable the model to achieve more reasonable results.
Multi-Task Learning. The diversity of anime styles presents a challenge for video generation. Although our model performs well in most styles, unique styles may result in inconsistencies, particularly in character details. To address this, we applied multi-task learning, combining image and video training to enhance the model’s adaptability to diverse styles.
We evaluated multi-task training using a manga with a unique artistic style. About 270 illustrations were used for the image generation task, while video training data remained the same as the baseline model. Additional illustrations served as first-frame conditions during video generation. After 5k training steps, as shown in Fig. 7, without incorporating images, the model struggles to fully understand such styles, resulting in flaws in character detail generation. While with the help of a small dataset of 270 images, the generated videos showed significantly greater stability and improved visual quality, particularly with highly distinctive guidance images. This approach effectively tailors animations to specific characters and mitigates domain gaps caused by variations in artistic styles, especially when high-quality animation data is limited.
Low-resolution vs High-resolution.
@@ 5.3 Animation Video Training
During the weak-to-strong training process, we observed that higher frame rates and resolutions enhance stability in visual details. As illustrated in Fig. 8, at 480P, facial features exhibit noticeable distortions, while at 720P, the model preserves both motion consistency and fine details. The higher resolution increases token representation for high-density areas, improving temporal consistency and overall content quality.
@@
6 Conclusion
In this paper, our proposed AniSora, a unified framework provides a solution to overcoming the challenges in animation video generation. Our data processing pipeline generates over 10M high-quality training clips, providing a solid base for our model. Leveraging a spatiotemporal mask, the generation model can create videos based on diverse control conditions. Furthermore, our evaluation benchmark demonstrates the effectiveness of our method in terms of character consistency and motion smoothness. We hope that our research and evaluation dataset establish a new benchmark and inspire further work in the animation industry.
Despite the promising results, some artifacts and flickering issues are still present in our generated animation videos. In the future, we aim to develop a comprehensive automated scoring system specifically designed for animation video evaluation datasets, ensuring closer alignment with human subjective perceptions. Additionally, we plan to expand the current model architecture to incorporate guidance across multiple modalities, such as camera movements, trajectories, skeletal motions, and audio. To tackle the challenge posed by the limited availability of high-quality animation data, we will employ reinforcement learning techniques to further refine the model’s performance.

### v4

@@ 6 Experiment
Table 1: Benchmark Evaluation Results
Models Human Visual Visual Visual Text-Video Image-Video Character
Evaluation Smooth Motion Appeal Consistency Consistency Consistency
Vidu-1.5 60.98 55.37 78.95 50.68 60.71 66.85 82.57
Opensora-V1.2 41.10 22.28 74.9 22.62 52.19 55.67 74.76
Opensora-Plan-V1.3 46.14 35.08 77.47 36.14 56.19 59.42 81.19
CogVideoX-5B-V1 53.29 39.91 73.07 39.59 67.98 65.49 83.07
MiniMax-I2V01 69.63 69.38 68.05 70.34 76.14 78.74 89.47
AniSora(Ours) 70.13 71.47 47.94 64.44 72.92 81.54 94.54
AniSora(Interpolated Average) - 70.78 53.02 64.41 73.56 80.62 91.59
AniSora(KeyFrame Interpolation) - 70.03 58.1 64.57 74.57 80.78 91.98
Figure 5: Human Evaluation and Benchmark Results Alignment
@@ 6.1 Benchmark Evaluation
In this section, we give both objective and human evaluation results of our benchmark. Six recent i2v models are involved in our evaluation: Open-sora Zheng et al. (2024), Open-sora-plan Lab and etc. (2024), Cogvideox Yang et al. (2024), Vidu-1.5 Vidu , Minimax-I2V01 Minimax and Anisora(ours). Tab. 1 gives the detailed scores from $6$ dimensions in the benchmark evaluation and the overall scores of the human evaluation. We observe that our model performs better than the other five methods on most dimensions, especially on visual smoothness and character consistency, except on the visual motion dimension. These mainly because we conduct a thorough assessment of the balance between generation quality and motion magnitude, and find most generation clips with big motion results in distortion or unnatural segments. In order to prove our benchmark is applicable to anime scenarios, we also evaluated our benchmark dataset using seven relevant dimensions in VBench benchmark. Due to the space limitation, we outline the results in Appendix. We observed that certain dimensions, including Motion Smoothness, Aesthetic Quality, I2V Background, and Overall Consistency, lacked sufficient discriminative power. In particular, some poorly generated results received higher scores than the ground truth, highlighting a discrepancy that fails to accurately capture human perception and experience.
Fig. 5 illustrates the detailed correlations among $6$ dimensions between human evaluation and benchmark results. Obviously, they are highly correlated with each other. Furthermore, we will publish our benchmark soon.
@@ 6.2 Spatiotemporal Mask Module
Frame Interpolation Tab. 1 presents the results of different interpolation settings on benchmark dataset. Our evaluation process involved generating videos on our benchmark with various guidance conditions sampled at equal proportions, which can refer to Fig .3. We then compute the average score of all samples as well as a specific statistical analysis for keyframe interpolation results. The performance indicates that single-frame guidance achieves competitive results whether the guiding frame is placed at the beginning, middle, or end of the frame sequence, which also consistently outperforms other methods. Adding more guiding frames further improves character consistency. We also observed from the motion and smooth score that our baseline model achieves a balance between motion range and consistency, while keyframe guidance enables the model to produce animation videos with larger motion ranges and more realistic motion. More samples can be found in the appendix.
Motion Area Condition The evaluation of motion area condition is constructed based on our benchmark dataset. For each initial frame, we performed saliency segmentation, followed by connected-component analysis to generate bounding boxes for each instance. Then we manually filtered the results to select high-quality motion area masks, resulting in 200 samples. Following the experiment settings in Dai et al. (2023), we conducted the comparison of motion mask precision in Tab. 2. We also computed the score of AnimateAnything on our selected 200 samples. The lower score is primarily due to flickering and noise appearing outside the motion mask area. The results demonstrate the effectiveness of our spatial mask module in controlling movable regions. It is also noticeable that even without motion control, our generation model trained for animation video still shows a certain level of control. This may due to the effective prompt-based guidance for the main subject. Motion mask guidance examples are shown in the appendix.
@@
6.3 Animation Video Training
2D and 3D Animation Analysis using QWEN2 Wang et al. (2024) shows that 2D samples account for 85% of our data set, yet the quality of 3D animation generation consistently exceeds that of 2D. Benchmark evaluations in the appendix confirm 3D animations demonstrate superior visual appearance and consistency, a phenomenon unique to animation training. We attribute this gap to the pre-trained model’s exposure to real-world video data. Unlike 2D animations with diverse motion patterns, 3D animations rendered by physics-based engines like Unreal Engine follow consistent physical laws, enabling better knowledge transfer during SFT. Consequently, improving generalization on 2D animation data remains more challenging than on 3D or real-world data.
Multi-Task Learning We evaluated multi-task training using a manga with a unique artistic style. About 270 illustrations were used for the image generation task, while video training data remained the same as the baseline model. Additional illustrations served as first-frame conditions during video generation. After 5k training steps, as shown in the appendix, the generated videos showed significantly greater stability and improved visual quality, particularly with highly distinctive guidance images. This approach effectively tailors animations to specific characters and mitigates domain gaps caused by variations in artistic styles, especially when high-quality animation data is limited.
Low-resolution vs High-resolution During the weak-to-strong training process, we observed that higher frame rates and resolutions enhance stability in visual details. As demonstrated in the appendix, at 480P, facial features exhibit noticeable distortions, while at 720P, the model preserves both motion consistency and fine details. The higher resolution increases token representation for high-density areas, improving temporal consistency and overall content quality.
7 Conclusion
In this paper, our proposed AniSora, a unified framework provides a solution to overcoming the challenges in animation video generation. Our data processing pipeline generates over 10M high-quality training clips, providing a solid base for our model. Leveraging a spatiotemporal mask, the generation model can create videos based on diverse control conditions. Furthermore, our evaluation benchmark demonstrates the effectiveness of our method in terms of character consistency and motion smoothness. We hope that our research and evaluation dataset establish a new benchmark and inspire further work in the animation industry. Besides, we are going to evaluate more models on our benchmark, providing valuable insights for model optimization.
Despite promising results, some artifacts and flickering are still present in the generated videos. In future work, we plan to integrate reinforcement learning with our evaluation benchmark to generate higher-quality videos.
Acknowledgments
We sincerely thank Qiang Zhang, Xinwen Zhang, and Xingyu Zheng for their valuable support and insightful suggestions on improving the manuscript.
Supplementary Material
S1 Dataset and Benchmark
In this section, we give a description of the detail points in the Dataset section and Benchmark section in the body paper.
S1.1 Dataset
First, we give the distribution for dimensions (text-cover region/OCR score, optical flow score, aesthetic score, and number of frames) of the animation dataset for model training. (shown in Fig. 6)
Figure 6: Distribution of Duration, OCR, Aesthetics, and Optical Flow in the Animation Dataset for Training
As mentioned in the Animation Dataset Construction section on page 3, a few higher-quality clips will be finally filtered from the training set to further improve the model’s performance. Compared with the initial screening, we use the fine-tuned filter model and set more stringent thresholds from these four dimensions to select higher-quality clips. The count of remaining high-quality clips is $0.5\%$ of the total. Besides, we train several fine-grained data recognition models, e.g., talking, motion amplitude, and 2D/3D animation clip recognition. These models help us to control the distribution of the training clips, which can further improve the generation performance.
S1.2 Benchmark
Visual Motion When we train a motion model to evaluate the motion amplitude of the video clips (mentioned on page 5), the prompts in the inference step list as follows:
[’The protagonist has a large range of movement, such as running, jumping, dancing, or waving arms.’, ’The protagonist remains stationary in the video with no apparent movement.’ ]
We can get a $1\times 2$ similarity score from each input video clip and prompt. After the softmax function, we can get the final motion score.
Character Consistency Fig. 7 gives the framework of character consistency model on page 6. We leverage detection (GroudingDino), segmentation (SAM) and tracking tools to get the characters’ masks on ground truth clips at first. Then, we finetune a character encoder module to extractor corresponding characters’ features. In the infer step, we extract the characters’ features from generation video clips, and get the final score through similarity calculation.
Figure 7: The Framework of Character Consistency: Training and Inference
Figure 8: Comparision of 2D and 3D animation examples. The badcases in 2D animation are mainly due to exaggerated deformations, diverse appearances, and motions that violate physical laws.
Table 3: Quantitative Results on Vbench. (Note that AniSora-K denotes the results with keyframe interpolation, and AniSora-I denotes the interpolated average results of AniSora)
Method Appearance Consistency
Motion Aesthetic Imaging I2V I2V Overall Subject
Smoothness Quality Quality Subject Background Consistency Consistency
Opensora-Plan(V1.3) 99.13 53.21 65.11 93.53 94.71 21.67 88.86
Opensora(V1.2) 98.78 54.30 68.44 93.15 91.09 22.68 87.71
Vidu(V1.5) 97.71 53.68 69.23 92.25 93.06 20.87 88.27
Cogvideo(5B-V1) 97.67 54.87 68.16 90.68 91.79 21.87 90.29
MiniMax(I2V01) 99.20 54.66 71.67 95.95 95.42 21.82 93.62
AniSora 99.34 54.31 70.58 97.52 95.04 21.15 96.99
AniSora-K 99.12 53.76 68.68 95.13 93.36 21.13 94.61
AniSora-I 99.31 54.67 68.98 94.16 92.38 20.47 95.75
GT 98.72 52.70 70.50 96.02 95.03 21.29 94.37
Figure 10: Comparision of results w/wo multi-task learning. The highlighted regions in red demonstrate significant improvements in stability and consistency after applying multi-task learning.
S2 Method
S2.1 Supervised Fine-Tuning
More effective training strategies are listed as follows:
Weak to Strong Our video generation model adopts a weak-to-strong training strategy to progressively enhance its learning capabilities across varying resolutions and frame rates. Initially, the model is trained on 480P videos at 8fps for 3 epochs, allowing it to capture basic spatiotemporal dynamics at a lower frame rate. Following this, the model undergoes training on 480P videos at 16fps for an additional 1.9 epochs, enabling it to refine its temporal consistency and adapt to higher frame rates. Finally, the model is fine-tuned on 720P videos at 16fps for 2.3 epochs, leveraging the previously learned features to generate high-resolution, temporally coherent video outputs. Additionally, we applied stricter filtering and produced a 1M ultra high-quality dataset for final-stage fine-tuning, significantly boosting high-resolution video quality.
Removing Generated Subtitles The presence of a significant number of videos with subtitles and platform watermarks in our training data led to the model occasionally generating such artifacts in its outputs. To mitigate this issue, we performed supervised fine-tuning using a curated dataset of videos entirely free of subtitles and watermarks. This dataset, consisting of 790k video clips, was constructed through proportional cropping of videos containing subtitles and the selection of clean, subtitle-free videos. Full-parameter fine-tuning was then applied to the model, and after 5.5k iterations, we observed that the model effectively eliminated the generation of subtitles and watermarks without compromising its overall performance.
Temporal Multi-Resolution Training Given the scarcity of high-quality animation data, we employ a mixed training strategy using video clips of varying durations to maximize data utilization. Specifically, a variable-length training approach is adopted, with training durations ranging from 2 to 8 seconds. This strategy enables our model to generate 720p video clips with flexible lengths between 2 and 8 seconds.
@@ Acknowledgments
S3 Experiment
S3.1 Quantitative Results on Vbench
Tab. 3 gives the results on $7$ dimensions from Vbench. It is easy to observe that the scores between different methods are a lack of differentiation.
Table 4: Benchmark Evaluation Results between 2D and 3D Generation Clips
Dims 3D 2D All
Visual Smooth 69.44 71.69 71.47
Visual Motion 45.59 70.29 47.94
Visual Appeal 57.99 65.04 64.44
Text-Video Consistency 72.86 72.93 72.92
Image-Video Consistency 79.75 82.29 81.54
Character Consistency 93.5 94.64 94.54
S3.2 2D and 3D Animation
The Fig. 8 demonstrates some results of 2D and 3D animation generation. As we discussed in the Experiment section, despite having a larger number of 2D data, generating high-quality 2D animation videos remains challenging due to the complexity of animation. Artifacts are more prevalent in 2D generation results, such as exaggerated deformations, more diverse character appearances, and motions that break the physical rules. For instance, in the third row of 2D examples, tears appear to floating in the air, making it more difficult for the model to capture the dynamic details accurately. In contrast, the motions rendered by the physics-based engines in 3D animations enable the model to achieve more reasonable results.
Moreover, Tab. 4 gives the benchmark results among 2D, 3D, and total clips. The results support the observation that 3D generation results perform better than those of 2D, especially on motion score.
S3.3 Spatiotemporal Mask Module
S3.3.1 Keyframe Guide Examples
@@
S3.3.2 Low-resolution vs High-resolution
Fig. 12 shows the comparison between 480P and 720P, where higher resolution improves temporal consistency and visual detail stability.
S3.3.3 Motion Area Condition
Precise and effective control is especially crucial in the animation production pipeline. Fig. 11 illustrates several motion mask guidance examples.
@@
S3.3.4 Multi-Task Learning
The diversity of anime styles presents a challenge for video generation. Although our model performs well in most styles, unique styles may result in inconsistencies, particularly in character details. To address this, we applied multi-task learning, combining image and video training to enhance the model’s adaptability to diverse styles. As shown in Fig. 10, without incorporating images, the model struggles to fully understand such styles, resulting in flaws in character detail generation. With the help of a small dataset of 270+ images, multi-task learning significantly improves the model’s ability to generate consistent and accurate details. With this technique, our unified framework can quickly adapt to more sophisticated styles, addressing the varied demands of creators.
High-resolution fine-tuning enhances the model’s ability to capture intricate image details, as illustrated in Fig. 12. At a resolution of 480p, distortions in facial features can be observed. However, at 720p, the model not only maintains dynamic consistency but also preserves fine details in areas such as faces and hands.
S4 More Examples
In this section, we show more generation examples compared with other methods, including Vidu, open-sora, open-sora plan, cogvideox, and minimax. Our unified framework can effectively generate higher-quality animation videos that align more closely with the prompt while avoiding issues such as character distortion or motion blur. Examples are shown in Fig. 13.